# Unearthing Gems from Stones: Policy Optimization with Negative Sample Augmentation for LLM Reasoning

### v3

@@ Abstract
Zhaohui Yang1,3thanks: Equal contribution., Yuxiao Ye2,3∗, Shilei Jiang3, Chen Hu3, Linjing Li1, Shihong Deng3thanks: Corresponding author. Work is done at StepFun., Daxin Jiang3 1 Institute of Automation, Chinese Academy of Sciences 2 Beijing Institute of Technology 3StepFun, China
@@ 1 Introduction
Reasoning capabilities are a critical aspect of evaluating the intelligence of large language models (LLMs). Recent advances have witnessed a paradigm shift from short to long chain-of-thought (CoT) reasoning, particularly after the release of OpenAI’s o1 series OpenAI (2024) and Deepseek R1 Guo et al. (2025). However, generating lengthy CoT responses incurs substantial computational costs. In the online learning paradigm, the periodic rollout process has become the bottleneck of post-training system Luo et al. (2025), significantly reducing training efficiency. Consequently, maximizing the utility of fixed training datasets under offline learning paradigm becomes increasingly crucial.
@@
Experimental results demonstrate that BCPG-NSA achieves the best performance and improved sample efficiency compared to baselines on challenging math reasoning benchmarks AIME24 and AIME25 and o.o.d coding benchmark LiveCodeBench. In addition, our ablation studies show that the consensus-based LLM-PRM annotation approach yields the best results compared to LLM-only and PRM-only approaches, indicating the importance of precise step correctness assessment in negative samples. We also demonstrate that BCPG-NSA is robust across different mining coefficient values, maintains stability over extended training epochs, and scales effectively to multiple iterations.
@@ 2 Related Work
Benefits from Negative Data  Many RL methods for LLMs, such as GRPO Shao et al. (2024) and GPG Chu et al. (2025), consider every step in a negative sample incorrect, and use the same strength to push down the likelihood of all tokens in incorrect responses. Prior research on learning from negative samples primarily focus on training data construction. Recent works on DPO Rafailov et al. (2023); Setlur et al. (2024) proposes constructing training pairs with shared prefixes between positive and negative samples, aiming to improve the model’s decision-making at critical intermediate steps. Another research targeting SFT data construction Wang et al. (2024) additionally add a prefix to indicate whether the current generation is a successful trajectory, helping the model better distinguish between correct and incorrect responses. The most closely related work Li et al. (2024b) leverages valuable signals from negative samples in short CoT reasoning, through a specialized dual-LoRA framework for model distillation. Under the long CoT paradigm, incorrect and correct steps more frequently alternate throughout the reasoning process. Therefore, fine-grained mining of the value of negative samples is a promising direction.
@@ 3 Preliminary Analysis: The Value of Negative Samples
Figure 1: (a): A case study: presence of correct steps (via human judgments) within incorrect response. (b): The overall framework of BCPG-NSA.
@@ 4.1 Segmenting Thinking Process into Steps
where ||||| | is the concatenation operator, and ${\rm start}_{k}$ and ${\rm end}_{k}$ represent the indices of the starting and ending tokens of the $k$-th step, respectively.
@@ 4.2 Consensus-based Annotation by LLM and PRM Judgers
After performing segmentation on the negative samples, we use a LLM judger and a discriminative PRM to jointly annotate each step in the reasoning process as correct/incorrect.
@@ 5.1.1 Training Dataset Construction
For the offline RL training dataset, we build the prompt seed set based on open-source data collected by the Open-Reasoner-Zero project Hu et al. (2025), comprising AIME (up to 2023), OpenR1-Math-220k Ben Allal et al. (2025) and various other open-source datasets Li et al. (2024a); Lambert et al. (2025); Hendrycks et al. (2021). During data filtering, we exclude multiple-choice and true/false questions to prevent cases where the model might occasionally derive the correct answer through incorrect reasoning steps, which could result in false positives and introduce noise into the training process.
@@ 5.1.4 Baselines
  • GRPO-offline Shao et al. (2024): It applies the GRPO loss consistently throughout the offline RL training process, eliminates the need for periodic online data resampling.
@@ 6.1 Different Annotation Methods
Results are demonstrated in Table 5. Compared to the base model DS-R1-14B and the vanilla BCPG, negative sample augmentation consistently improves performance across all 3 annotation methods, demonstrating the robustness of NSA’s benefits regardless of the choice of judgers. Moreover, the LLM-PRM approach, despite mining the smallest number of correct tokens (only 26M in total) from negative samples, achieves the best performance. We hypothesize that the consensus-based filtering implements more stringent selection criteria, ensuring higher quality of retained correct steps and avoiding potential false positives from individual judgers.
@@ 6.2 Different Values of Mining Coefficient
Mining coefficient $\beta$ is the key hyperparameter in NSA, used to control the strength of the augmentation. Therefore, we test the performance of the model under different $\beta$ values within the range $[-1,1]$ to demonstrate the robustness of NSA. When $\beta=1$, BCPG-NSA reduces to the vanilla BCPG method.
@@ 8 Conclusion
In this paper, we present BCPG-NSA, a novel offline RL framework designed to effectively leverage valuable components within negative samples for LLMs’ long CoT reasoning. Unlike existing methods that either discard negative samples or apply equal penalization, our approach enables fine-grained treatment of negative samples through three key components: semantic step segmentation, consensus-based step correctness assessment, and policy optimization with negative sample augmentation. Experimental results on challenging reasoning benchmarks demonstrate that BCPG-NSA achieves superior performance, and exhibits great robustness and scalability when extended to multiple iterations.
@@ Appendix A Case Study: A negative response that incorporates positive steps
Problem: What is the sum of all of the multiples of 7 between 100 and 200? Solution: Step 1 Content: Hmm, okay. Let me think about this step by step. Maybe I can start by recalling how to find the sum of an arithmetic series because multiples of 7 should form an arithmetic sequence. Wait, arithmetic series? Yes, because each term increases by a constant difference, which in this case is 7. So, I can use the formula for the sum of an arithmetic series. … Then, dividing both sides by 7, I get 13 = n - 1. Adding 1 to both sides, n = 14. So, there are 14 terms in this sequence. Score: 1. Reason: Correctly identifies the first term (105), last term (196), and number of terms (14) in the arithmetic sequence. Step 2 Content: Wait, let me verify that. Let’s list out some terms to be sure. Starting at 105: 105, 112, 119, 126, 133, 140, 147, 154, 161, 168, 175, 182, 189, 196. Yeah, that’s 14 terms. Okay, so n = 14. Now, plugging these values into the sum formula: $S=n/2*(\mathrm{first\ term}+\mathrm{last\ term})$. So, that would be $S=14/2*(105+196)$. Calculating that, 14 divided by 2 is 7, and 105 plus 196 is 299. So, 7 multiplied by 299.Hmm, 7*300 is 2100, so 7*299 is 2100 minus 7, which is 2093.Wait, really? … Score: 0. Reason: Miscalculates the sum by incorrectly adding the first and last terms (105 + 196 = 299 instead of 301) and incorrectly computes the sum up to 200 using the wrong first term (105 instead of 7). Step 3 Content: Wait, that’s a different number. So, which one is correct? … Alternatively, since all these numbers form an arithmetic progression, I can use the formula again. So, $S=n/2*(\mathrm{first\ term}+\mathrm{last\ term})$. So, n is 14, first is 105, last is 196. So, 14/2 is 7, times (105 + 196) is 7*299=2093. … Score: 0. Reason: Continues based on STEP 2’s errors (299 instead of 301 and incorrect sum up to 200) without correction or a new correct approach. Step 4 Content: … Wait, maybe I miscalculated one of the sums. Let me compute the sum from 105 to 196 manually, but that would be time-consuming, but let me try a few terms. 105 + 112 + 119 + 126 + 133 + 140 + 147 + 154 + 161 + 168 + 175 + 182 + 189 + 196. Let me add them step by step. Start with 105. 105 + 112 = 217. 217 + 119 = 336. 336 + 126 = 462. 462 + 133 = 595. 595 + 140 = 735. 735 + 147 = 882. 882 + 154 = 1036. 1036 + 161 = 1197. 1197 + 168 = 1365. 1365 + 175 = 1540. 1540 + 182 = 1722. 1722 + 189 = 1911. 1911 + 196 = 2107. Score: 1. Reason: Correctly verifies the sum as 2107 through manual addition and the difference method, despite repeating an earlier incorrect calculation.
@@
[…]
@@ Appendix B Prompt Template for LLM Judger
You are an expert in evaluating mathematical problem-solving processes. The user will provide a math problem, a step-by-step solution and ground truth. Your tasks are as follows: 1. First, extract a concise final answer (Short Answer) from the ground truth. 2. Then, carefully review the user’s step-by-step solution, assigning a score to each step (either 0 or 1). For each step, provide a brief explanation of your judgement result. Scoring rules: If a step contains an explicit error, such as a reasoning error, calculation mistake, logical flaw, or misunderstanding of the problem, it should be scored 0. If a step does not contain any errors, score it according to the following rules: 1. Error Propagation: If a previous step contains an error and the current step continues the analysis based on that error without introducing a new, correct approach or making a proper correction, the current step should also be scored 0. 2. Error Termination: If a previous step contains an error, but the current step either corrects the previous mistake or introduces a new and correct approach, the current step should be scored 1. For example: STEP K contains an error. STEP K+1 continues the analysis based on the error. STEP K+2 corrects the previous error or introduces a new and correct approach. In this case, STEP K and STEP K+1 should be scored 0, and STEP K+2 should be scored 1. Your response format should be in json format:
@@
Figure 6: Prompt template for LLM judger.
@@ Appendix C Training Details
$J$ (${}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\bigg{[}$
$\log\sigma\left(\beta_{\rm DPO}\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.$ (8)
$\left.\left.-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right)\bigg{]}$

### v4

@@ Abstract
Zhaohui Yang1,3thanks: Equal contribution., Yuxiao Ye2,311footnotemark: 1, Shilei Jiang3, Chen Hu3, Shihong Deng3thanks: Corresponding author., Linjing Li1, Daxin Jiang3, 1Institute of Automation, Chinese Academy of Sciences 2Beijing Institute of Technology  3StepFun, China yangzhaohui2023@ia.ac.cnyuxiaoye99cs@gmail.com dshnightmare@gmail.com
@@ 1 Introduction
Reasoning capabilities are a critical aspect of evaluating the intelligence of large language models (LLMs). Recent advances have witnessed a paradigm shift from short to long chain-of-thought (CoT) reasoning, particularly after the release of OpenAI’s o1 series OpenAI (2024) and Deepseek R1 Guo et al. (2025). However, generating lengthy CoT responses incurs substantial computational costs. In the online learning paradigm, the periodic rollout process has become the bottleneck of the post-training system Luo et al. (2025), significantly reducing training efficiency. Consequently, maximizing the utility of fixed training datasets under offline learning paradigm becomes increasingly crucial.
@@
Experimental results demonstrate that BCPG-NSA achieves the best performance and improved sample efficiency compared to baselines on challenging math reasoning benchmarks AIME24 and AIME25, and o.o.d coding benchmark LiveCodeBench. In addition, our ablation studies show that the consensus-based LLM-PRM annotation approach yields the best results compared to LLM-only and PRM-only approaches, indicating the importance of precise step correctness assessment in negative samples. We also demonstrate that BCPG-NSA is robust across different mining coefficient values, maintains stability over extended training epochs, and scales effectively to multiple iterations.
@@ 2 Related Work
Benefits from Negative Data  Many RL methods for LLMs, such as GRPO Shao et al. (2024) and GPG Chu et al. (2025), consider every step in a negative sample incorrect, and use the same strength to push down the likelihood of all tokens in incorrect responses. Prior research on learning from negative samples primarily focuses on training data construction. Recent works on DPO Rafailov et al. (2023); Setlur et al. (2024) propose constructing training pairs with shared prefixes between positive and negative samples, aiming to improve the model’s decision-making at critical intermediate steps. Another research targeting SFT data construction Wang et al. (2024) additionally adds a prefix to indicate whether the current generation is a successful trajectory, helping the model better distinguish between correct and incorrect responses. The most closely related work Li et al. (2024b) leverages valuable signals from negative samples in short CoT reasoning, through a specialized dual-LoRA framework for model distillation. Under the long CoT paradigm, incorrect and correct steps more frequently alternate throughout the reasoning process. Therefore, fine-grained mining of the value of negative samples is a promising direction.
@@ 3 Preliminary Analysis: The Value of Negative Samples
Figure 1: (a): A case study: presence of correct steps (via human judgments) within an incorrect response. (b): The overall framework of BCPG-NSA.
@@ 4.1 Segmenting Thinking Process into Steps
where $||$ is the concatenation operator, and ${\rm start}_{k}$ and ${\rm end}_{k}$ represent the indices of the starting and ending tokens of the $k$-th step, respectively.
@@ 4.2 Consensus-based Annotation by LLM and PRM Judgers
After performing segmentation on the negative samples, we use an LLM judger and a discriminative PRM to jointly annotate each step in the reasoning process as correct/incorrect.
@@ 5.1.1 Training Dataset Construction
For the offline RL training dataset, we build the prompt seed set based on open-source data collected by the Open-Reasoner-Zero project Hu et al. (2025), comprising AIME (up to 2023), OpenR1-Math-220k Ben Allal et al. (2025), and various other open-source datasets Li et al. (2024a); Lambert et al. (2025); Hendrycks et al. (2021). During data filtering, we exclude multiple-choice and true/false questions to prevent cases where the model might occasionally derive the correct answer through incorrect reasoning steps, which could result in false positives and introduce noise into the training process.
@@ 5.1.4 Baselines
  • GRPO-offline Shao et al. (2024): It applies the GRPO loss consistently throughout the offline RL training process, eliminating the need for periodic online data resampling.
@@ 6.1 Different Annotation Methods
Results are demonstrated in Table 5. Compared to the base model DS-R1-14B and the vanilla BCPG, negative sample augmentation consistently improves performance across all 3 annotation methods, demonstrating the robustness of NSA’s benefits regardless of the choice of judgers. Moreover, the LLM-PRM approach, despite mining the smallest number of correct tokens (only 26M in total) from negative samples, achieves the best performance. We hypothesize that the consensus-based filtering implements more stringent selection criteria, ensuring a higher quality of retained correct steps and avoiding potential false positives from individual judgers.
@@ 6.2 Different Values of Mining Coefficient
Mining coefficient $\beta$ is the key hyperparameter in NSA, used to control the strength of the augmentation. Therefore, we test the performance of the model under different $\beta$ values within the range [1,1][-1,1] to demonstrate the robustness of NSA. When $\beta=1$, BCPG-NSA reduces to the vanilla BCPG method.
@@ 8 Conclusion
In this paper, we present BCPG-NSA, a novel offline RL framework designed to effectively leverage valuable components within negative samples for LLMs’ long CoT reasoning. Unlike existing methods that either discard negative samples or apply equal penalization, our approach enables fine-grained treatment of negative samples through three key components: semantic step segmentation, consensus-based step correctness assessment, and policy optimization with negative sample augmentation. Experimental results on challenging reasoning benchmarks demonstrate that BCPG-NSA achieves superior performance and exhibits great robustness and scalability when extended to multiple iterations.
@@ 9 Limitations
10 Acknowledgments
This work was supported in part by the Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDA0480301.
@@ Appendix A Case Study: A negative response that incorporates positive steps
Problem: What is the sum of all of the multiples of 7 between 100 and 200? Solution: Step 1 Content: Hmm, okay. Let me think about this step by step. Maybe I can start by recalling how to find the sum of an arithmetic series because multiples of 7 should form an arithmetic sequence. Wait, arithmetic series? Yes, because each term increases by a constant difference, which in this case is 7. So, I can use the formula for the sum of an arithmetic series. … Then, dividing both sides by 7, I get 13 = n - 1. Adding 1 to both sides, n = 14. So, there are 14 terms in this sequence. Score: 1. Reason: Correctly identifies the first term (105), last term (196), and number of terms (14) in the arithmetic sequence. Step 2 Content: Wait, let me verify that. Let’s list out some terms to be sure. Starting at 105: 105, 112, 119, 126, 133, 140, 147, 154, 161, 168, 175, 182, 189, 196. Yeah, that’s 14 terms. Okay, so n = 14. Now, plugging these values into the sum formula: $S=n/2*(\mathrm{first\ term}+\mathrm{last\ term})$. So, that would be S=14/2(105+196)S=14/2*(105+196). Calculating that, 14 divided by 2 is 7, and 105 plus 196 is 299. So, 7 multiplied by 299.Hmm, 7*300 is 2100, so 7*299 is 2100 minus 7, which is 2093.Wait, really? … Score: 0. Reason: Miscalculates the sum by incorrectly adding the first and last terms (105 + 196 = 299 instead of 301) and incorrectly computes the sum up to 200 using the wrong first term (105 instead of 7). Step 3 Content: Wait, that’s a different number. So, which one is correct? … Alternatively, since all these numbers form an arithmetic progression, I can use the formula again. So, $S=n/2*(\mathrm{first\ term}+\mathrm{last\ term})$. So, n is 14, first is 105, last is 196. So, 14/2 is 7, times (105 + 196) is 7*299=2093. … Score: 0. Reason: Continues based on STEP 2’s errors (299 instead of 301 and incorrect sum up to 200) without correction or a new correct approach. Step 4 Content: … Wait, maybe I miscalculated one of the sums. Let me compute the sum from 105 to 196 manually, but that would be time-consuming, but let me try a few terms. 105 + 112 + 119 + 126 + 133 + 140 + 147 + 154 + 161 + 168 + 175 + 182 + 189 + 196. Let me add them step by step. Start with 105. 105 + 112 = 217. 217 + 119 = 336. 336 + 126 = 462. 462 + 133 = 595. 595 + 140 = 735. 735 + 147 = 882. 882 + 154 = 1036. 1036 + 161 = 1197. 1197 + 168 = 1365. 1365 + 175 = 1540. 1540 + 182 = 1722. 1722 + 189 = 1911. 1911 + 196 = 2107. Score: 1. Reason: Correctly verifies the sum as 2107 through manual addition and the difference method, despite repeating an earlier incorrect calculation.
@@
Step 5 Content: …… Wait, so the manual addition gives me 2107, same as 2842 - 735.But when I used the arithmetic series formula, I got 2093. So, that’s the problem. Why is there a discrepancy? Wait, let me check the arithmetic series formula again. I think the issue is with the number of terms.Wait, the first term is 105, which is 15*7, and the last term is 196, which is 28*7. Score: 1. Reason: Correctly applies the alternative arithmetic series formula to get 2107, confirming the correct sum. Step 6 Content: Wait, earlier when I did the arithmetic series formula correctly, I think I messed up. Wait, no. If I use $S=n/2*(\mathrm{first}+\mathrm{last})$, which is $(14/2\times(105+196)=7\times 299=2093$. But when we I used the correct formula S=n/2(2a+(n1)d)S=n/2*(2a+(n-1)d), it gave me 2107. Wait, that’s a contradiction. So, what’s happening here. Wait, perhaps I made an error in the number of terms. … So, n=14, first term 105, last term 196. So, the sum is 14/2(105+196)=7299=209314/2*(105+196)=7*299=2093 Score: 0. Reason: Reintroduces the incorrect sum (2093) alongside the correct 2107 without resolving the error. Step 7-14. A series of deductions based on false premises. Step 15 Content: So, there are 14 terms in this sequence. We then use the arithmetic series sum formula: $S=\frac{n}{2}\times(\mathrm{first\ term}+\mathrm{last\ term})$. Substituting the values, we get: $S=\frac{14}{2}\times(105+196)=7\times 299=2093$. Thus, the sum of all multiples of 7 between 100 and 200 is $\boxed{2093}$. Score: 0. Reason: Concludes with the incorrect sum of 2093 instead of the correct 2107.
@@ Appendix B Prompt Template for LLM Judger
You are an expert in evaluating mathematical problem-solving processes. The user will provide a math problem, a step-by-step solution, and ground truth. Your tasks are as follows: 1. First, extract a concise final answer (Short Answer) from the ground truth. 2. Then, carefully review the user’s step-by-step solution, assigning a score to each step (either 0 or 1). For each step, provide a brief explanation of your judgment result. Scoring rules: If a step contains an explicit error, such as a reasoning error, calculation mistake, logical flaw, or misunderstanding of the problem, it should be scored 0. If a step does not contain any errors, score it according to the following rules: 1. Error Propagation: If a previous step contains an error and the current step continues the analysis based on that error without introducing a new, correct approach or making a proper correction, the current step should also be scored 0. 2. Error Termination: If a previous step contains an error, but the current step either corrects the previous mistake or introduces a new and correct approach, the current step should be scored 1. For example: STEP K contains an error. STEP K+1 continues the analysis based on the error. STEP K+2 corrects the previous error or introduces a new and correct approach. In this case, STEP K and STEP K+1 should be scored 0, and STEP K+2 should be scored 1. Your response format should be in json format:
@@
Figure 6: Prompt template for the LLM judger.
@@ Appendix C Training Details
$J$ $_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\bigg{[}$
[…]